what is the value of j in the equation

/j + /j + 14 = 3 /j + 10 ?

Answer by checkley71(8403)

1/J+1/J+14=3/J+10 OR 2/J+14=3/J+10 OR 14-10=3/J-2/J OR 4=1/J OR J=1/4
PROOF 1/(1/4)+1/(1/4)+14=3/(1/4)+10 OR 4+4+14=12+10 OR 22=22